OpenStudy (anonymous):

Can someone please demonstrate how to turn a radical expression into an exponent expression? And then how to turn an exponent expression into a radical expression? Please and thanks so much!

OpenStudy (anonymous):

I know how to simplify a radical expression, but that is about it. I am completely lost here, and have to answer this question for school today, so if someone could please help me I would greatly appreciate it. Thanks.

OpenStudy (kirbykirby):

\[\Large \left(\sqrt[m]{x}\right)^n=\sqrt[m]{x^n}= x^{\frac{n}{m}}\]

OpenStudy (kirbykirby):

example: \[\Large \left(\sqrt[3]{125}\right)^7=\sqrt[3]{125^7}= 125^{\frac{7}{3}}\]

OpenStudy (kirbykirby):

In the "special" case where there is no number "in" the square root: \[\Large \sqrt{x}=x^{\frac{1}{2}} \]

OpenStudy (anonymous):

So if that is how to turn a radical expression into an exponent expression, then how do we turn an exponent expression into a radical expression?

OpenStudy (kirbykirby):

It's the same, just go in reverse. \[\Large x^{\frac{n}{m}}=\sqrt[m]{x^n}=\left(\sqrt[m]{x}\right)^n \]

OpenStudy (anonymous):

Thank you!